• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name X. Zheng

    Log entry time 16:00:40 on April30,2007

    Entry number 200636

    keyword=day shift summary

    11:38 Ron noticed ESR 4K flow dropped to zero, then is slowly back;
    11:40 lost right Q3, then all other magnets;
    12:30 cryo group has to shut off the flow to our target. The target vented;
    15:00 target is back, set the magnets to the correct momentum and waiting...
    16:00 right dipole is still not back, but won't take too long.

    Good production runs on 4He during this shift:

    3014-3023